OpenStudy (anonymous):

3x + 3y =12 5y+6x=7 find x and y

OpenStudy (anonymous):

\[3x + 3y = 12\]\[x+y = 4\]\[x =- y + 4\]\[5y + 6( - y + 4 ) = 7\]\[5y - 6y + 24 = 7\]\[-y +24 = 7\]\[-y = - 17\]\[y = 17\]

OpenStudy (anonymous):

\[3x + 3(17) = 12\]\[3x + 51 = 12\]\[3x = -39\]\[x = -13\]